Insert parentheses to make statement true:
30 - 4 x 2 + 5 = 2

Answer by ankor@dixie-net.com(22740)

Insert parentheses to make statement true:
30 - 4 x 2 + 5 = 2
:
30 - 4 x (2 + 5) = 2
:
30 - 4(7) = 2